Beth wants to make 11.6 fl. oz. of a 45% acid solution by mixing together a 22% acid solution and a 80% acid solution. How much of each solution must she use?
Answer by mananth(16946)

Beth wants to make 11.6 fl. oz. of a 45% acid solution by mixing together a 22% acid solution and a 80% acid solution. How much of each solution must she use?

Let the 22% solution required be x
The total required = 11.6 oz
22%------------------80%-----------45% strength
x--------------------11.6-x---------11.6 Quantity
0.22x+0.8(11-x) = 0.45*11.6
0.22x+8.8-0.8x=5.22
-0.58x==-3.58
x=-3.58/-0.58
= 6.17 oz of 22% acid

11.6-6.17 = 5.43 oz